A thief steals an ATM card and randomly must guess the correct five- digit pin code from a 6- key keypad. The number of possible codes is? The probability that the correct code is given on the first try is?
Answer #1

from a 6 digit key pad we can technically make

6*6*6*6*6 codes of 5 digit as all the 5 numbers of password can have 6 choices

Hence  The number of possible codes is = 6^5 =7776

So there are 7776 possible number of pin codes

Now  The probability that the correct code is given on the first try is =1/7776 = 0.0001286008
As there is only 1 possible correct ans outof 7776 possible combinations hence 0.0001286008 is the  probability that the correct code is given on the first try.

Hope the above answer has helped you in understanding the problem. Please upvote the ans if it has really helped you. Good Luck!!